A put option that expires in three months with an exercise price of $50 sells for $4.89. The stock is currently priced at $53, and the risk-free rate is 4.8 percent per year, compounded continuously. What is the price of a call option with the same exercise price? (Answer up to two decimal places)
Answer #1

Rate =4.8%
Time =3/12 =0.25 years
Call value - Stock Price + Strike Price/e^(rt) = Put Value
4.89-53+50/e^(4.8%*0.25) =Put Value
Price of a call option Put Value =1.29